A 4.65 kg cat is creeping up on a bird on a fence using a plank leaning on the fence. The plank makes a 42.5 º angle with the ground. What are the perpendicular and parallel components of the weight of the cat?.
1
Expert's answer
2020-11-16T07:45:47-0500


The weight of the cat has the magnitude


"P = mg = 4.65kg\times 9.81m/s^2 = 45.6165 \space N"

The parallel component is


"P_x = P\sin\alpha = 45.6165\cdot \sin42.5\degree \approx 30.82\cdot N"

The perpendicular component is


"P_y = P\cos\alpha = 45.6165\cdot \cos42.5\degree \approx 33.63\cdot N"

Answer. Parallel: 30.82 N, perpendicular: 33.63 N.